OpenStudy (anonymous):

Which expression is irrational? A. 49 B. 81 C. 99 D. 400

OpenStudy (anonymous):

None?

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

cause 99 squared is 9.9498744

OpenStudy (anonymous):

where 49 squared is 7, 81 = 9, and 400 = 20

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

mind helping me? http://openstudy.com/study#/updates/53705c6de4b09cc1503cb2c8

OpenStudy (tkhunny):

Rational Numbers can be expressed as the Quotient of Two Integers. Irrational Numbers cannot! A. 49 = 49/1 - Check B. 81 = 81/1 - Check C. 99 = 99/1 - Check D. 400 = 400/1 - Check

OpenStudy (anonymous):

Ohhh okk

OpenStudy (anonymous):

NONE !!!